OpenStudy (anonymous):

Sobean meal is 14% protein; cornmeal is 7% protein. How many pounds of each should be mixed together in order to get 280lb mixture that is 9% protein? How many pounds of the cornmeal should be in the mixture? How many pounds of soybean should e in the mixture?

OpenStudy (anonymous):

let s=14% =0.14 let c=7% =0.07 s+c=280 c =280 -s 0.14s+0.07c= 280* 0.09 can you solve from here ?

OpenStudy (anonymous):

yes thank you